# On the asymptotic behavior of the spectral gap for discrete Schrödinger operators

### v1

@@ Abstract.
In this note we elaborate on the convergence of the spectral gap of a class of discrete Schrödinger operators defined on a path graph in the limit of infinite volume. We confirm and extend recent results on the convergence of the spectral gap to a larger class of potentials using entirely different methods. Most importantly, we also resolve a conjecture previously proposed in this context. This then yields new insights into the rate at which the spectral gap tends to zero as the volume increases.
2010 Mathematics Subject Classification:
@@ 1. Introduction
This paper is devoted to the investigation of the spectral gap of discrete Schrödinger operators. Here, spectral gap refers to the difference between the lowest two eigenvalues. At least in the continuous setting, related considerations have a long tradition: for example, important results were obtained in [AB89] and then in [Lav94], solving the one-dimensional version of the fundamental gap conjecture [AC11]. As a matter of fact, it turns out that it is in general quite difficult to say something about how the gap changes under perturbations, for instance, under the addition of some non-negative potential. As a matter of fact, it might already be difficult to infer the correct sign of the changes [Abr91]. This is certainly true as long as the configuration space on which the operator is defined remains fixed; for example, one could think of a Schrödinger operator defined on an interval of fixed length [AK20, ACH21]. However, as demonstrated recently [KT23, KT24], one is able to say more about the spectral gap for a larger class of potentials if one is working with Schrödinger operators defined on domains of increasing volume. Most importantly, in [KT24] the authors discussed an interesting effect which was recently studied on the path graph [KY25] and which forms – in some sense – the starting point of our investigations. More explicitly, it was shown that the presence of (certain) compactly supported potentials drastically change the asymptotic behavior of the spectral gap when compared to the asymptotic behavior of the gap of the Laplacian without a potential. In other words, as soon as one adds such a non-negative compactly supported potential to the Laplacian, the spectral gap converges strictly faster to zero in the limit of infinite volume. This somewhat surprising effect is the result of an effective degeneracy of the lowest two eigenvalues at infinite volume. It is important to note, however, that this effect might disappear in higher dimensions or reappear as well, depending on the particular nature of the potential [KT23]. This also explains why the authors of [KY25] focused on a specific discrete graph – the path graph: adding a compactly supported potential to the Laplacian then leads, at least on an intuitive level, to the mentioned (approximate) degeneracy of the lowest two eigenvalues in the infinite-volume limit since the potential cuts the graph into two (approximately) congruent pieces.
@@
Starting with the results of [KY25], we want to achieve two things with this paper: First, since the authors of [KY25] considered only a very specific potential localized on only one vertex, we generalize their main result [KY25, Theorem 6] to general non-negative and compactly supported potentials, while using entirely different proof techniques. Furthermore, we prove the conjecture put forward in [KY25, Section 4] which provides more information on how fast the spectral gap converges to zero. By doing this, we provide some rigorous justification to the numerical results obtained in [KY25, Section 4].
@@
For completeness, let us also mention that investigations of the spectral gap (or spectral properties) for Jacobi operators [Moe76, Tes00] and Schrödinger-type operators on graph-like structures such as combinatorial graphs [KLW21] and metric graphs [BK13, Kur24] are quite common. Regarding the discrete case, let us refer to [FCLP18, LS20, JMM22, BS24]; for the metric case, one may consult [KGMN13, KKMM16, BL17, Roh17, BKKM17] and references therein.
@@
Our paper is organized as follows: In Section 2 we introduce the basic setting and the class of Schrödinger operators considered. In Section 3 we then derive upper and lower bounds on the two lowest eigenvalues. Finally, in Section 4 we use these estimates to prove our main result (Theorem 7) and to resolve a conjecture from [KY25, Section 4] (Corollary 9).
@@ 2. The model
In the following, we study operators defined on the finite-dimensional Hilbert space $\mathcal{H}_{k}=\mathds{C}^{|V_{k}|}$, $k\in\mathds{N}$. The discrete (standard, unweighted) Laplacian $\mathcal{L}_{\gamma,k}:\mathds{C}^{|V_{k}|}\rightarrow\mathds{C}^{|V_{k}|}$ is defined via
@@
where $J\subset\mathds{Z}$ is a non-empty set independent of k; $\alpha_{j}>0$, $j\in J$, represents the strength of the external potential localized at the $j$-th vertex. The associated self-adjoint operator then becomes
@@
where $\boldsymbol{\alpha}=(\alpha_{j})_{j\in J}$ and $\delta_{j}$ is such that $(\delta_{j}f)_{i}:=\delta_{ij}f(j)$, $i\in V_{k}$, $\delta_{ij}$ referring to the Kronecker delta. As in the case for $\mathcal{L}_{k}$, the operator $H_{k,\boldsymbol{\alpha}}$ is, for all $k\in\mathds{N}$, non-negative and its spectrum consists of non-negative eigenvalues only. For all $k\in\mathds{N}$, we denote the eigenvalues of $H_{k,\boldsymbol{\alpha}}$ by
@@
Note that $\Gamma(V_{k},\boldsymbol{\alpha})>0$ for each value of $k\in\mathds{N}$, since the ground state is non-degenerate; more on the spectral theory of graphs can be found in [Chu97, BH12, KLW21]. For convenience, we set
$\widetilde{\alpha}\coloneq\min\limits_{j\in J}\alpha_{j}\qquad\text{ and }\qquad\widehat{\alpha}\coloneq\sum\limits_{j\in J}\alpha_{j}\ .$
It is important to mention that for the special case $J=\{0\}$ with external potential $\alpha\delta_{0}$, one has
for all $\alpha\geq 0$. Here, $\lambda_{1}(V_{k},\alpha\delta_{0})$ is shorthand for $\lambda_{1}(V_{k},\boldsymbol{\alpha})$ with external potential $\alpha\delta_{0}$ and $\lambda_{0}(V_{k},\infty\delta_{0})$ is shorthand for the lowest eigenvalue of the Laplacian on the path graph with a Dirichlet condition at the zero vertex. Later we shall also use that there exist constants $c_{1},c_{2}>0$ such that for all but finitely many $k\in\mathds{N}$,
$c_{1}k^{-2}\leq\lambda_{0}(V_{k},\infty\delta_{0})\leq c_{2}k^{-2}\ ;$ (5)
this follows from a well-known explicit formula for $\lambda_{1}(V_{k},\alpha\delta_{0})$ as used, for example, in [KY25]. In the following, we will also frequently use the identity
@@
We conclude this section with a remark to help the reader better understand our notation used in the following.
@@
Firstly, throughout this paper, we denote by $\varphi_{k}$ the normalized ground state of $H_{k,\boldsymbol{\alpha}}$, that is, the eigenvector corresponding to the lowest eigenvalue $\lambda_{0}(V_{k},\boldsymbol{\alpha})$ with $\|\varphi_{k}\|^{2}_{\mathcal{H}_{k}}=\sum_{j=-k}^{k}|\varphi_{k}(j)|^{2}=1$. Note that the ground state $\varphi_{k}$ can be chosen to be strictly positive.
Secondly, for two sequences $(a_{k})_{k\in\mathds{N}}$, $(b_{k})_{k\in\mathds{N}}$, $a_{k}\neq 0$ for all but finitely many $k\in\mathds{N}$, we use the notation $b_{k}\lesssim a_{k}$ iff there is a constant $c>0$ such that $|b_{k}/a_{k}|\leq c$ for all but finitely many $k\in\mathds{N}$, and $b_{k}\sim a_{k}$ iff there are constants $c_{1},c_{2}>0$ such that $c_{1}\leq|b_{k}/a_{k}|\leq c_{2}$ for all but finitely many $k\in\mathds{N}$.
Also, in the proofs we denote by $(const.)$ positive constants that are independent of $k$ and the potential, whenever their precise values are irrelevant to the core arguments of our proofs.
@@ 3. Spectral estimates
$a_{k,1}\coloneq 1/2-\sum_{j=-k}^{r_{\min}}|\varphi_{k}(j)-\varphi_{k}(r_{\min})|^{2}\quad$ $\text{ and }\quad a_{k,2}\coloneq 1/2-\sum_{j=r_{\max}}^{k}|\varphi_{k}(j)-\varphi_{k}(r_{\max})|^{2}\ .$
Note that $1\geq a_{k,1}+a_{k,2}\geq 1-\sum_{j=-k}^{k}|\varphi_{k}(j)|^{2}=0$ due to monotonicity of the ground state outside of $r_{\min},r_{\max}]$.
@@
$\lambda_{0}(V_{k},\boldsymbol{\alpha})$ $\geq\Big{(}\dfrac{1}{2}-a_{k,1}\Big{)}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})+\Big{(}\dfrac{1}{2}-a_{k,2}\Big{)}\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\ .$ (6)
@@
$0\leq a_{k,1}+a_{k,2}\lesssim\widetilde{\alpha}^{-1}k^{-1}\ .$ (7)
@@
In a first step we obtain
$\lambda_{0}(V_{k},\boldsymbol{\alpha})$ $\geq\sum\limits_{j=-k}^{r_{\min}-1}|\nabla_{j}(\varphi_{k}-\varphi_{k}(r_{\min}))|^{2}+\sum\limits_{j=r_{\max}}^{k}|\nabla_{j}(\varphi_{k}-\varphi_{k}(r_{\max}))|^{2}+\sum\limits_{j\in J}\alpha(j)|\varphi_{k}(j)|^{2}$
$\geq\sum\limits_{j=-k}^{r_{\min}}|\varphi_{k}(j)-\varphi_{k}(r_{\min})|^{2}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})$
$\qquad+\sum\limits_{j=r_{\max}}^{k}|\varphi_{k}(j)-\varphi_{k}(r_{\max})|^{2}\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})+\sum\limits_{j\in J}\alpha(j)|\varphi_{k}(j)|^{2}\ .$
Using the definition of $a_{k,1},a_{k,2}$ and the fact that the last term is non-negative then implies the lower bound (6).
@@
$\sum\limits_{j\in J}\alpha(j)|\varphi_{k}(j)|^{2}\leq a_{k,1}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})+a_{k,2}\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\ ,$ (8)
@@
$\sum\limits_{j\in J}|\varphi_{k}(j)|^{2}\leq\widetilde{\alpha}^{-1}\Big{(}a_{k,1}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})+a_{k,2}\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\Big{)}\ .$ (9)
@@
$1=\sum\limits_{j=-k}^{k}|\varphi_{k}(j)|^{2}=\sum\limits_{j=-k}^{r_{\min}-1}|\varphi_{k}(j)|^{2}+\sum\limits_{j=r_{\max}+1}^{k}|\varphi_{k}(j)|^{2}+\sum\limits_{j=r_{\min}}^{r_{\max}}|\varphi_{k}(j)|^{2}$
@@
$\begin{split}a_{k,1}+a_{k,2}&=\sum\limits_{j=r_{\min}}^{r_{\max}}|\varphi_{k}(j)|^{2}+2\varphi_{k}(r_{\min})\sum\limits_{j=-k}^{r_{\min}-1}\varphi_{k}(j)+2\varphi_{k}(r_{\max})\sum\limits_{j=r_{\max}+1}^{k}\varphi_{k}(j)\\ &\qquad\qquad-\varphi_{k}^{2}(r_{\min})(k+r_{\min})-\varphi_{k}^{2}(r_{\max})(k-r_{\max})\ .\end{split}$ (10)
@@
$a_{k,1}+a_{k,2}$ $\leq(const.)\max\Big{\{}\widetilde{\alpha}^{-1/2}\left(a_{k,1}+a_{k,2}\right)^{1/2}\Theta_{\max}^{1/2}k^{1/2},\sum\limits_{j=r_{\min}}^{r_{\max}}|\varphi_{k}(j)|^{2}\Big{\}}\ .$
In order to estimate the last sum, we use that for all $j\in\{r_{\min},\ldots,r_{\max}\}$, again due to (9) and the eigenvalue equation,
$|\varphi_{k}(j)|^{2}$ $\leq\left(|\varphi_{k}(r_{\min})|+\sum_{i=r_{\min}}^{j-1}|\varphi_{k}(i+1)-\varphi_{k}(i)|\right)^{2}$ (11)
$\leq 2\max_{j\in J}|\varphi_{k}(j)|^{2}+2|r_{\max}-r_{\min}|\lambda_{0}(V_{k},\alpha)\ .$
Therefore, using $a+b\leq\max\{2a,2b\}$ for $a,b>0$, we get
$\begin{split}a_{k,1}+a_{k,2}&\leq(const.)\max\Big{\{}\widetilde{\alpha}^{-1}\Theta_{\max}k,\Theta_{\max}\Big{\}}\ .\end{split}$ (12)
The claimed estimate (7) now follows by observing that $\Theta_{\max}\sim k^{-2}$.
@@
Inspecting the proof of Theorem 2, in particular (8), we obtain the following estimate for the special case when $J=\{0\}$, that is, for $H_{k,\alpha}=\mathcal{L}_{k}+\alpha\delta_{0}$ with $\alpha>0$: There exists a constant $c>0$ such that, for all but finitely many $k\in\mathds{N}$,
$\varphi_{k}(0)\leq c\alpha^{-1}k^{-3/2}\ .$
This can be regarded as a generalization of [KY25, Lemma 5]. Note here that $\varphi_{k}(0)=(2k+1)^{-1/2}$ for $\alpha=0$.
From a physics point of view, it is interesting to remark that the just mentioned estimate implies an estimate on the potential energy of a particle described by $H_{k,\alpha}=\mathcal{L}_{k}+\alpha\delta_{0}$ with $\alpha>0$: More explicitly, the potential energy in the ground state is given by $E_{pot}(\alpha):=\alpha|\varphi_{k}(0)|^{2}$ and we therefore obtain, for $k\in\mathds{N}$,
$E_{pot}(\alpha)\leq c^{2}\alpha^{-1}k^{-3}$
with the constant $c>0$ from above. The interesting point here is that the potential energy is of a lower order than the kinetic energy, which is of order $\sim k^{-2}$.
Remark 4.
We encourage the reader to keep the following physical intuition in mind, which also served as a guiding perspective in the development of this work: Our focus lies on the two lowest eigenvalues, both of which converge to zero as the size of the graph tends to infinity. In this regime, a particle occupying a quantum state associated with either of these eigenvalues has arbitrarily low energy. As a result, any potential of finite strength effectively appears to be of an infinite height to such a low-energy particle. As a consequence, this provides some explanation for the emergence of quantities in our statements related to infinitely strong potentials.
@@
$\begin{split}\widetilde{\cos}_{k,r_{\min}}(j)&\coloneq\begin{cases}A_{\min}^{-1/2}\cos\left(\left[j+k+\frac{1}{2}\right]\pi/(2(k+r_{\min})+1)\right),\ -k\leq j\leq r_{\min}\ ,\\ 0\ \quad\text{else}\ ,\end{cases}\\ \widetilde{\cos}_{k,r_{\max}}(j)&\coloneq\begin{cases}A_{\max}^{-1/2}\cos\left(\left[k-j+\frac{1}{2}\right]\pi/(2(k-r_{\max})+1)\right),\ r_{\max}\leq j\leq k\ ,\\ 0\ \quad\text{else}\ .\end{cases}\end{split}$
Here, $A_{\max},A_{\min}>0$ are such that both functions are normalized to $1/2$. Also note that both functions are constructed using the well-known explicit form of the second eigenfunction to the Hamiltonian $H_{k,\boldsymbol{\alpha}}=\mathcal{L}_{k}+\alpha\delta_{0}$.
Theorem 5 (Upper bound ground state energy).
Consider a Hamiltonian $H_{k,\boldsymbol{\alpha}}$ with ground state $\varphi_{k}\in\mathcal{H}_{k}$ corresponding to the ground state energy $\lambda_{0}(V_{k},\boldsymbol{\alpha})$. Then

### v2

@@ Abstract.
In this note we elaborate on the asymptotic behavior of the spectral gap of a class of discrete Schrödinger operators defined on a path graph in the limit of infinite volume. We confirm recent results and generalize them to a larger class of potentials using entirely different methods. Notably, we also resolve a conjecture previously proposed in this context. This then yields new insights into the rate at which the spectral gap tends to zero as the volume increases.
M. Hofmann was supported by the Fundação para a Ciência e a Tecnologia (FCT), Portugal, within the scope of the projects Spectral Optimal Partitions: geometric and numerical analysis, reference 2023.13921.PEX.
@@ 1. Introduction
This paper is devoted to the investigation of the spectral gap of discrete Schrödinger operators. Here, the spectral gap refers to the difference between the lowest two eigenvalues. At least in the continuous setting, related considerations have a long tradition: for example, important results were obtained in [AB89] and then in [Lav94], solving the one-dimensional version of the fundamental gap conjecture [AC11]. As a matter of fact, it turns out that it is in general quite difficult to say something about how the gap changes under perturbations, for instance, under the addition of some non-negative potential. Indeed, it might already be difficult to infer the correct sign of the changes [Abr91]. This is certainly true as long as the configuration space on which the operator is defined remains fixed; for example, one could think of a Schrödinger operator defined on an interval of fixed length [AK20, ACH21]. However, as demonstrated recently [KT23, KT24], one is able to say more about the spectral gap for a larger class of potentials if one is working with Schrödinger operators defined on domains of increasing volume. Most importantly, in [KT24] the authors discussed an interesting effect which was recently studied on the path graph [KY25] and which forms – in some sense – the starting point of our investigations. More explicitly, it was shown that the presence of (certain) compactly supported potentials drastically change the asymptotic behavior of the spectral gap when compared to the asymptotic behavior of the gap of the Laplacian without a potential. In other words, as soon as one adds such a non-negative compactly supported potential to the Laplacian, the spectral gap converges strictly faster to zero in the limit of infinite volume. This somewhat surprising effect is the result of an effective degeneracy of the lowest two eigenvalues at infinite volume. It is important to note, however, that this effect might disappear in higher dimensions or reappear as well, depending on the particular nature of the potential [KT23]. This also explains why the authors of [KY25] focused on a specific discrete graph – the path graph: adding a compactly supported potential to the Laplacian then leads, at least on an intuitive level, to the mentioned (approximate) degeneracy of the lowest two eigenvalues in the infinite-volume limit since the potential cuts the graph into two (approximately) congruent pieces.
@@
Starting with the results of [KY25], we want to achieve two things with this paper: First, since the authors of [KY25] considered only a very specific potential localized on only one vertex, we generalize their main result [KY25, Theorem 6] to general non-negative and compactly supported potentials, while using entirely different proof techniques. Furthermore, we prove the conjecture put forward in [KY25, Section 4], which provides more information on how fast the spectral gap converges to zero. By doing this, we also provide some rigorous justification to the numerical results obtained in [KY25, Section 4].
@@
For completeness, let us also mention that investigations of the spectral gap (or spectral properties) for Jacobi operators [Moe76, Tes00] and Schrödinger-type operators on graph-like structures such as combinatorial graphs [KLW21] and metric graphs [BK13, Kur24] are quite common. Regarding the discrete case, let us refer to [FCLP18, LS20, JMM22, BS24]; for the metric case, one may consult [KMN13, KKMM16, BL17, Roh17, BKKM17] and references therein.
@@
Our paper is organized as follows: In Section 2 we introduce the basic setting and the class of Schrödinger operators considered. In Section 3 we then derive upper and lower bounds on the two lowest eigenvalues. Finally, in Sections 4 and 5, we use these estimates to prove our main results (Theorem 7 and Theorem 11) and hereby resolve a conjecture from [KY25, Section 4], see in particular Corollary 9 and Theorem 11. In the Appendix A, we derive auxiliary results that we use in Section 5.
@@ 2. The model
In the following, we study operators defined on the finite-dimensional Hilbert space $\mathcal{H}_{k}=\mathds{C}^{|V_{k}|}$, $k\in\mathds{N}$. The discrete (standard, unweighted) Laplacian $\mathcal{L}_{k}:\mathds{C}^{|V_{k}|}\rightarrow\mathds{C}^{|V_{k}|}$ is defined via
@@
where $\boldsymbol{\alpha}=(\alpha_{j})_{j\in J}$ and $J\subset\mathds{Z}$ is a non-empty set independent of $k$; $\alpha_{j}>0$, $j\in J$, represents the strength of the external potential localized at the $j$-th vertex. The associated self-adjoint operator then becomes
@@
where $\delta_{j}$ is such that $(\delta_{j}f)_{i}:=\delta_{ij}f(j)$, $i\in V_{k}$, $\delta_{ij}$ referring to the Kronecker delta. As in the case for $\mathcal{L}_{k}$, the operator $H_{k,\boldsymbol{\alpha}}$ is, for all $k\in\mathds{N}$, non-negative and its spectrum consists of non-negative eigenvalues only. For all $k\in\mathds{N}$, we denote the eigenvalues of $H_{k,\boldsymbol{\alpha}}$ by
@@
Note that $\Gamma(V_{k},\boldsymbol{\alpha})>0$ for each value of $k\in\mathds{N}$, since the ground state is non-degenerate; more on the spectral theory of graphs can be found in [Chu97, BH12, KLW21]. We denote by $\varphi_{k,\boldsymbol{\alpha}}$ the normalized ground state of $H_{k,\boldsymbol{\alpha}}$, that is, the eigenvector corresponding to the lowest eigenvalue $\lambda_{0}(V_{k},\boldsymbol{\alpha})$ with $\|\varphi_{k,\boldsymbol{\alpha}}\|^{2}_{\mathcal{H}_{k}}=\sum_{j=-k}^{k}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}=1$. Note that the ground state $\varphi_{k,\boldsymbol{\alpha}}$ can be chosen to be strictly positive.
We set
$\widetilde{\alpha}\coloneq\min\limits_{j\in J}\alpha_{j}\qquad\text{ and }\qquad\widehat{\alpha}\coloneq\sum\limits_{j\in J}\alpha_{j}$
and frequently use the identity, $k\in\mathds{N}$,
@@
For the special case $J=\{0\}$ with external potential $\alpha\delta_{0}$, $\alpha>0$, we write
$H_{k,\alpha}\coloneq\mathcal{L}_{k}+\alpha\delta_{0}\ ,\quad k\in\mathds{N}\ ,$
for the Hamiltonian, $\varphi_{k,\alpha}$ for its ground state, and $\Gamma(V_{k},\alpha)$ for the spectral gap. It is important to mention that in this case, one has
for all $\alpha\geq 0$ and all $k\in\mathds{N}$. Here, since the external potential is $\alpha\delta_{0}$, we write $\lambda_{1}(V_{k},\alpha\delta_{0})$ for $\lambda_{1}(V_{k},\boldsymbol{\alpha})$; similarly, $\lambda_{0}(V_{k},\infty\delta_{0})$ is shorthand for the lowest eigenvalue of the Laplacian on the path graph with a Dirichlet condition at the zero vertex. In the same spirit, we use $\lambda_{0}(V_{k},\alpha\delta_{0})$ as shorthand for $\lambda_{0}(V_{k},\boldsymbol{\alpha})$ in case of an external potential $\alpha\delta_{0}$. Later we shall also use that, see for example [BH12],
$\lambda_{0}(V_{k},\infty\delta_{0})=2-2\cos\left(\pi/|V_{k}|\right)\ ,\quad k\in\mathds{N}\ ,$ (5)
thus, in particular,
$\lambda_{0}(V_{k},\infty\delta_{0})=\pi^{2}|V_{k}|^{-2}+\mathcal{O}(|V_{k}|^{-4})\ .$ (6)
We conclude this section with two remarks. The first is intended to clarify our notation used in this paper, while the second presents a physical intuition that motivated and guided our analysis.
@@
For two sequences $(a_{k})_{k\in\mathds{N}}$, $(b_{k})_{k\in\mathds{N}}$, $a_{k}\neq 0$ for all but finitely many $k\in\mathds{N}$, we use the notation $b_{k}\lesssim a_{k}$ iff there is a constant $c>0$ such that $|b_{k}/a_{k}|\leq c$ for all but finitely many $k\in\mathds{N}$, and $b_{k}\sim a_{k}$ iff there are constants $c,C>0$ such that $c\leq|b_{k}/a_{k}|\leq C$ for all but finitely many $k\in\mathds{N}$.
Also, in the proofs and some statements we denote by $(const.)$ positive constants that are independent of $k$ and the potential strength $\boldsymbol{\alpha}=(\alpha_{j})_{j\in J}$, whenever their precise values are irrelevant. Note, however, that they could still depend on the support of the potential.
Remark 2.
We encourage the reader to keep the following physical intuition in mind, which also guided the development of this work. Our analysis focuses on the two lowest eigenvalues, both of which converge to zero as the size of the graph tends to infinity. In this low-energy regime, a particle occupying a quantum state corresponding to either eigenvalue has arbitrarily small kinetic energy. Consequently, any potential of fixed finite strength, informally, becomes effectively impenetrable to such a particle, and therefore behaves, in the limit, as if it were infinitely strong. This perspective helps to explain the appearance of quantities related to infinitely strong potentials in our results.
@@ 3. Spectral estimates
$a_{k,1}$ $\coloneq\dfrac{1}{2}-\sum_{j=-k}^{r_{\min}}|\varphi_{k,\boldsymbol{\alpha}}(j)-\varphi_{k,\boldsymbol{\alpha}}(r_{\min})|^{2}\ ,$ (7)
as well as
$a_{k,2}$ $\coloneq\dfrac{1}{2}-\sum_{j=r_{\max}}^{k}|\varphi_{k,\boldsymbol{\alpha}}(j)-\varphi_{k,\boldsymbol{\alpha}}(r_{\max})|^{2}\ .$ (8)
Note that $1\geq a_{k,1}+a_{k,2}\geq 1-\sum_{j=-k}^{k}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}=0$ due to monotonicity of the ground state outside of $r_{\min},r_{\max}]$.
@@
$\begin{split}\lambda_{0}(V_{k},\boldsymbol{\alpha})&\geq\left(\dfrac{1}{2}-a_{k,1}\right)\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})+\left(\dfrac{1}{2}-a_{k,2}\right)\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\\ &\qquad+\sum\limits_{j\in J}\alpha_{j}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}\ .\end{split}$ (9)
@@
$0\leq a_{k,1}+a_{k,2}\leq(const.)\widetilde{\alpha}^{-1}k^{-1}$ (10)
for all but finitely many $k\in\mathds{N}$.
@@
In a first step we obtain, using the minmax-principle,
$\lambda_{0}(V_{k},\boldsymbol{\alpha})$ $\geq\sum\limits_{j=-k}^{r_{\min}-1}|\nabla_{j}(\varphi_{k,\boldsymbol{\alpha}}-\varphi_{k,\boldsymbol{\alpha}}(r_{\min}))|^{2}+\sum\limits_{j=r_{\max}}^{k}|\nabla_{j}(\varphi_{k,\boldsymbol{\alpha}}-\varphi_{k,\boldsymbol{\alpha}}(r_{\max}))|^{2}$
$\qquad+\sum\limits_{j\in J}\alpha_{j}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}$
$\geq\sum\limits_{j=-k}^{r_{\min}}|\varphi_{k,\boldsymbol{\alpha}}(j)-\varphi_{k,\boldsymbol{\alpha}}(r_{\min})|^{2}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})$
$\qquad+\sum\limits_{j=r_{\max}}^{k}|\varphi_{k,\boldsymbol{\alpha}}(j)-\varphi_{k,\boldsymbol{\alpha}}(r_{\max})|^{2}\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})+\sum\limits_{j\in J}\alpha_{j}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}\ .$
Using the definition of $a_{k,1},a_{k,2}$ then implies the lower bound (9).
@@
$\sum\limits_{j\in J}\alpha_{j}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}\leq a_{k,1}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})+a_{k,2}\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\ ,$ (11)
@@
$\sum\limits_{j\in J}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}\leq\widetilde{\alpha}^{-1}\Big(a_{k,1}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})+a_{k,2}\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\Big)\ .$ (12)
@@
$1=\sum\limits_{j=-k}^{k}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}=\sum\limits_{j=-k}^{r_{\min}-1}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}+\sum\limits_{j=r_{\max}+1}^{k}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}+\sum\limits_{j=r_{\min}}^{r_{\max}}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}$
@@
$\begin{split}a_{k,1}+a_{k,2}&=\sum\limits_{j=r_{\min}}^{r_{\max}}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}+2\varphi_{k,\boldsymbol{\alpha}}(r_{\min})\sum\limits_{j=-k}^{r_{\min}-1}\varphi_{k,\boldsymbol{\alpha}}(j)+2\varphi_{k,\boldsymbol{\alpha}}(r_{\max})\sum\limits_{j=r_{\max}+1}^{k}\varphi_{k,\boldsymbol{\alpha}}(j)\\ &\qquad\qquad-|\varphi_{k,\boldsymbol{\alpha}}(r_{\min})|^{2}(k+r_{\min})-|\varphi_{k,\boldsymbol{\alpha}}(r_{\max})|^{2}(k-r_{\max})\ .\end{split}$ (13)
@@
$a_{k,1}+a_{k,2}$ $\leq(const.)\max\Big\{\widetilde{\alpha}^{-1/2}\left(a_{k,1}+a_{k,2}\right)^{1/2}\Theta_{\max}^{1/2}k^{1/2},\sum\limits_{j=r_{\min}}^{r_{\max}}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}\Big\}\ .$
In order to estimate the last sum, we use that for all $j\in\{r_{\min},\ldots,r_{\max}\}$, since $(a+b)^{2}\leq 2a^{2}+2b^{2}$ for $a,b\in\mathds{R}$ and due to the eigenvalue equation on the form level,
$|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}$ $\leq\left(|\varphi_{k,\boldsymbol{\alpha}}(r_{\min})|+\sum_{i=r_{\min}}^{j-1}|\varphi_{k,\boldsymbol{\alpha}}(i+1)-\varphi_{k,\boldsymbol{\alpha}}(i)|\right)^{2}$ (14)
$\leq 2\max_{j\in J}|\varphi_{k,\boldsymbol{\alpha}}(j)|^{2}+2|r_{\max}-r_{\min}|\lambda_{0}(V_{k},\boldsymbol{\alpha})\ .$
Therefore, using again the inequalities $a+b\leq\max\{2a,2b\}$ for $a,b>0$ and (12), we get
$\begin{split}a_{k,1}+a_{k,2}&\leq(const.)\max\Big\{\widetilde{\alpha}^{-1}\Theta_{\max}k,\Theta_{\max}\Big\}\end{split}$ (15)
for all but finitely many $k\in\mathds{N}$. The claimed estimate (10) now follows by observing that $\Theta_{\max}\sim k^{-2}$.
@@
Inspecting the proof of Theorem 3, in particular (11), we obtain the following estimate for the special case when $J=\{0\}$, that is, for $H_{k,\alpha}=\mathcal{L}_{k}+\alpha\delta_{0}$ with $\alpha>0$: For all but finitely many $k\in\mathds{N}$,
$\varphi_{k,\alpha}(0)\leq(const.)\alpha^{-1}k^{-3/2}\ .$ (16)
This can be regarded as a generalization of [KY25, Lemma 5]. Note here that $\varphi_{k,0}(0)=|V_{k}|^{-1/2}$ (that is for $\alpha=0$).
From a physics point of view, it is interesting to remark that the just mentioned estimate implies an upper bound on the potential energy of a particle described by $H_{k,\alpha}=\mathcal{L}_{k}+\alpha\delta_{0}$ with $\alpha>0$: More explicitly, the potential energy in the ground state is given by $E_{pot}(\alpha):=\alpha|\varphi_{k,\alpha}(0)|^{2}$ and we therefore obtain, for $k\in\mathds{N}$,
$E_{pot}(\alpha)\leq(const.)\alpha^{-1}k^{-3}\ .$
The interesting point here is that the potential energy is of a lower order than the kinetic energy, which is of order $k^{-2}$.
@@
$\begin{split}\widetilde{\cos}_{k,r_{\min}}^{\text{l}}(j)&\coloneq\begin{cases}A_{\min}^{-1/2}\cos\left(\left(j+k+\frac{1}{2}\right)\pi/(2(k+r_{\min})+1)\right),\ -k\leq j\leq r_{\min}\ ,\\ 0\ \quad\text{else}\ ,\end{cases}\\ \widetilde{\cos}_{k,r_{\max}}^{\text{r}}(j)&\coloneq\begin{cases}A_{\max}^{-1/2}\cos\left(\left(k-j+\frac{1}{2}\right)\pi/(2(k-r_{\max})+1)\right),\ r_{\max}\leq j\leq k\ ,\\ 0\ \quad\text{else}\ .\end{cases}\end{split}$ (17)
Here, $A_{\max},A_{\min}>0$ are such that both functions are normalized to $1/2$. Also note that both functions are constructed using the well-known explicit form of the second eigenfunction to the Hamiltonian $H_{k,\boldsymbol{\alpha}}=\mathcal{L}_{k}+\alpha\delta_{0}$. In particular, $\varphi_{k,\infty}=\widetilde{\cos}_{k,0}^{\text{l}}+\widetilde{\cos}_{k,0}^{\text{r}}$ where $\varphi_{k,\infty}$ denotes the ground state of $H_{k,\alpha}$ for $\alpha=\infty$.
Theorem 5 (Upper bound ground state energy I).
Consider a Hamiltonian $H_{k,\boldsymbol{\alpha}}$ with ground state $\varphi_{k,\boldsymbol{\alpha}}\in\mathcal{H}_{k}$ corresponding to the ground-state energy $\lambda_{0}(V_{k},\boldsymbol{\alpha})$. Then